Please help my solve this. I have been trying to figure it out.
Find three equations that have an x-intercept at (1,0)

Answer by Alan3354(69443)

Find three equations that have an x-intercept at (1,0)
-------------
x = 1
x = y + 1
x = -y + 1
etc
x = ky + 1, k = any value